# Prompt-based Dynamic Token Pruning to Guide Transformer Attention in Efficient Segmentation

### v1

@@ Abstract
The high computational demands of Vision Transformers (ViTs), in processing a huge number of tokens, often constrain their practical application in analyzing medical images. This research proposes an adaptive prompt-guided pruning method to selectively reduce the processing of irrelevant tokens in the segmentation pipeline. The prompt-based spatial prior helps to rank the tokens according to their relevance. Tokens with low-relevance scores are down-weighted, ensuring that only the relevant ones are propagated for processing across subsequent stages. This data-driven pruning strategy facilitates end-to-end training, maintains gradient flow, and improves segmentation accuracy by focusing computational resources on essential regions. The proposed framework is integrated with several state-of-the-art models to facilitate the elimination of irrelevant tokens; thereby, enhancing computational efficiency while preserving segmentation accuracy. The experimental results show a reduction of $\sim$ 35-55% tokens; thus reducing the computational costs relative to the baselines. Cost-effective medical image processing, using our framework, facilitates real-time diagnosis by expanding its applicability in resource-constrained environments.
Keywords:
@@ 1 Introduction
Precise segmentation of anatomical or pathological areas in medical images is important for clinical applications, such as disease diagnosis, treatment planning, surgical navigation, and personalized medicine. It allows clinicians to measure the burden of the disease, evaluate the response to treatment, and ultimately improve the patient’s outcome. Deep learning [1] has revolutionized medical image analysis by offering powerful algorithms to automate complex tasks such as segmentation. Its ability to learn intricate patterns from large datasets without much human intervention has made it a popular choice. Vision Transformers (ViTs) [2] have recently gained prominence in the development of algorithms for various computer vision tasks, including medical image segmentation.
@@
ViTs employ self-attention mechanism to capture long-range dependencies in images, thus facilitating global context awareness. This is beneficial for segmentation of anatomical structures that exhibit diverse shapes and sizes. The global view enables ViTs to model interactions between far-apart image regions, which can be difficult for convolutional networks with their restricted receptive fields. Several ViT-based segmentation approaches viz. UNETR [3], TransUNET [4], Swin-UNETR [5] and SegFormer [6] have reported encouraging performance in medical image segmentation, achieving state-of-the-art results across various benchmarks.
@@
Despite impressive performance, ViTs face limitation in handling resource-constrained environments due to the quadratic computational complexity of self-attention. This scales poorly with increasing input image dimensions. Consequently, the deployment of ViT-based segmentation models in real-time clinical environments becomes challenging, limiting their applicability in practice.
@@
Model compression techniques, including pruning, offer an effective way to mitigate the computational expenses associated with ViTs. Pruning reduces the computational cost of a model by removing less significant elements from its architecture. Static pruning [7, 8, 9] removes fixed components of the model architecture, before inference, by irreversibly eliminating components considered less significant. Although effective in minimizing model size and computational expense, static pruning is less flexible [10]. Its rigid nature limits the ability to adjust to the unique features of each input image. This constraint poses significant challenges for the analysis of medical images, where the variability between images can be considerable.
In contrast, dynamic pruning [11] adaptively adjusts the computation at inference by selectively deactivating model components based on input. Token pruning increases efficiency by reducing the number of processed tokens. Early exit strategies [12, 13] save computation by bypassing the processing of specific layers for a subset of tokens. However, they might introduce representation inconsistencies within the network. The importance scores for weighting tokens in dynamic token pruning approaches [14, 15, 16] are often learned without explicit consideration of prior spatial knowledge about the varying target anatomical structures. This might lead to suboptimal segmentation results.
@@
To address these limitations, we introduce a Prompt-driven Adaptive Token Pruning approach that explicitly incorporates structured spatial priors into the pruning process. In contrast to techniques that depend exclusively on attention scores or heuristics for selecting tokens, our method utilizes the spatial information embedded in the box prompts to direct token retention. It facilitates input-specific adaptation of the network computation. This guarantees the preservation of tokens essential for depicting the structure of the target object, while redundant or less informative tokens are removed. By directly incorporating this spatial constraint, the proposed method improves the precision of the segmentation and optimizes computational efficiency.
@@
Section 2 provides a comprehensive overview of the steps involved in the proposed token pruning framework. The experimental results and details of implementation are presented in Section 3, in the publicly available datasets ACDC [17] and ISIC [18]. Qualitative and quantitative results derived from the application of our framework, as embedded in state-of-the-art ViT-based segmentation models, are described. Section 4 provides a summary of the concluding remarks.
2 Methodology
This section details our proposed Prompt-Driven Token Pruning framework, in the context of the architecture of prevalent segmentation networks, as well as the token generation mechanism adopted in ViTs. Recent high-performance medical image segmentation models employ an $U-$ shaped encoder-decoder architecture [19]. ViT-based models typically employ ViT blocks either at the bottleneck of the encoder-decoder structure or throughout the entire encoder path. This study focuses on architectures associated with both frameworks.
2.1 Token generation in ViT
@@
ViTs transform an input image (or feature map volume) $I\in\mathbb{R}^{C\times H\times W}$ into a sequence of tokens [2]. Here, $H$, $W$ and $C$ represent the height, width, and channel dimensions of $I$. The input is partitioned into a set of non-overlapping patches $Z;|Z|=\frac{HW}{p^{2}}$, with each patch having dimension $p^{2}$. They are flattened and linearly projected in $C^{\prime}$-dimensional embedding space, resulting in the sequence $P^{\prime}\in\mathbb{R}^{Z\times C^{\prime}}$. These embedded vectors are called tokens. Positional embedding is added to the tokens to retain their spatial locations. $P_{0}\in\mathbb{R}^{Z\times C^{\prime}}$ is the sequence of position-sensitive tokens that serve as input to a transformer encoder.
@@
This encoder block consists of alternating layers of multi-head self-attention ($\mu$) and feed-forward network ($\Phi$) units. The $\mu$ captures global contextual relationships between different tokens, performing self-attention parallelly across multiple heads, to operate in various representational subspaces. $\Phi$ is independently applied to the output of each head, to enhance their representational capacity by including an additional non-linear transformation. The operations within a transformer block are expressed as $Y_{out}=\Phi(LN(Y))+\mu(LN(P_{0}))+P_{0}$, where $LN$ [20] is the Layer Normalization and $Y_{out}\in\mathbb{R}^{Z\times C^{\prime}}$ is the final output of the transformer block.
2.2 Prompt-driven adaptive pruning
@@
Prompts serve as a prior by offering high-level, task-specific information. This directs the model to focus on segmenting specific structures within the input image, resulting in effective segmentation. Consequently, it allows token selection based on contextual information. This framework is referred to as prompt-driven adaptive pruning, where essential tokens are preserved while down-weighting the irrelevant ones to minimize their processing.
Score map compute: At a given stage $l$, employing a ViT block, the output tokens $Y_{out}$ are combined with box-prompt coordinates $(x_{min},y_{min}),(x_{max},y_{max})$ to extract region-specific features $F\in\mathbb{R}^{k\times k\times C^{\prime}}$ from the area defined by the prompt using RoIAlign [21]. The box prompts serve as spatial priors, guiding the selection of relevant features by focusing on the most informative regions. This facilitates the extraction of region-specific feature embeddings corresponding to the Region-of-Interest (RoI) defined in the prompt.
@@
In addition, a similarity score map $S\in\mathbb{R}^{M\times Z}$ is calculated between the region-specific feature, encoded by $F$ and $Y_{out}$, to measure the relevance of each token. The $F$ and $Y_{out}$ are first transformed into lower-dimensional embeddings $V_{1}\in\mathbb{R}^{M\times d_{V}}$ and $V_{2}\in\mathbb{R}^{Z\times d_{V}}$, respectively, with projection matrices $f_{1}\in\mathbb{R}^{C^{\prime}\times d_{V}}$ and $f_{2}\in\mathbb{R}^{C^{\prime}\times d_{V}}$. This reduces computational complexity while improving feature discrimination. $M$ denotes the number of feature vectors in $V_{1}$. The similarity score is computed as
$S(V_{1},V_{2})=V_{1}V_{2}^{T}/\sqrt{d_{V}};V_{1}=f_{1}.F;V_{2}=f_{2}.Y_{out}.$ (1)
@@
Prune mask compute: Here $S$ is normalized along the token dimension to generate a probability distribution $\rho\in\mathbb{R}^{M\times Z}$. The probability $\rho_{ij}$ of the $i$ th feature vector of $V_{1}$ ($V_{1i}$) that it is relevant to the $j$th feature vector of $V_{2}$ ($V_{2j}$) is given as $\rho_{ij}=\frac{e^{(S_{ij})}}{\sum_{k=1}^{Z}e^{S_{ik}}}$. $S_{ij}$ denotes the similarity score between $V_{1i}$ and $V_{2j}$. Next, the Shannon entropy [22] $H(.)$ is calculated for the probability distribution of each $V_{1i}$, over all the ViT tokens. This yields
@@
with $H(V_{1}{{}_{i}})$ quantifying the uncertainty of association between $V_{1i}$ and the set of all ViT tokens. A lower value of $H(V_{1}{{}_{i}})$ implies a strong association with a subset of tokens. Conversely, a higher value indicates that $V_{1i}$ is uncertain about its association with the tokens.
An inverse entropy weighting scheme is applied to prioritize tokens having lower uncertainty values. The entropy values are first ranked in ascending order. The ranks $R_{i}|i\in[1,M]$ are normalized to ensure their uniform spacing between 0 and 1. Inverse entropy weights $\Omega_{i}=1-R_{i}$, with $\Omega=\{\Omega_{i}|i\in[1,M]\}$, ensure that higher entropy values (more uncertainty) are assigned to the lower weights and vice versa.
Lemma 1 (Uniformity of Weights)
@@
The weighted similarity score map becomes $\tilde{S}=S.\Omega$, with the overall relevance $r_{i}$ of the $i$th token being the mean of the weighted similarity scores
@@
The final step modifies the original tokens, based on their relevance scores. A binary mask $\beta\in\{0,1\}^{Z}$ is generated by applying a threshold $\tau$ on the relevance scores as
@@
Here $\sigma$ is the sigmoid function to squash the relevance scores to range $[0,1]$, and $\beta$ is applied to ViT tokens $Y_{out}$ for effectively removing tokens that are deemed irrelevant based on our proposed prompt-guided weighting and thresholding. The masked token $Y_{out}^{\prime}$ is generated by element-wise multiplication ($\odot$) of $\beta$ with $Y_{out}$, i.e. $Y_{out}^{\prime}=\beta\odot Y_{out}.$
3 Experimental Results
The proposed framework is evaluated on state-of-the-art ViT-based segmentation models, viz. UNETR, TransUNET, SegFormer and Swin UNETR. The Dice Score Coefficient (DSC) is used to compare the segmentation performance between the original baselines and their adaptations after incorporating our prompt-based token pruning. The evaluation is performed on publicly available ACDC [17] and ISIC [18] datasets. ACDC comprises 150 chest MRI images with ground truth delineations for right ventricle, left ventricle and myocardium. ISIC contains 2000 dermoscopic images with corresponding pixel-level annotations for skin lesions. The adopted train-test-validation split is 80:10:10. The models were trained in Python 3.9, using PyTorch and MONAI libraries, on NVIDIA RTX A5500 GPU with 24GB of memory.
Table 1: Ablation study of UNETR on ACDC dataset, over various threshold $\tau$.
$\tau$ = 0.3 0.5 Percentile
$25^{th}$ $50^{th}$ $75^{th}$
GFLOPs 13.32 9.07 26.78 37.31 18.66
mDSC 0.670 $\pm$ 0.07 0.661 $\pm$ 0.09 0.693 $\pm$ 0.07 0.695 $\pm$ 0.08 0.657 $\pm$ 0.09
Table 2: Ablation study of UNETR on ACDC dataset, with pruning performed at different stages.
Stage # {1} {1,2} {1,2,3} {1,2,3,4} Baseline
GFLOPs 37.42 26.78 37.35 37.31 37.46

### v2

# Prompt-based Dynamic Token Pruning for Efficient Segmentation of Medical Images

@@ Abstract
The high computational demands of Vision Transformers (ViTs) in processing a large number of tokens often constrain their practical application in analyzing medical images. This research proposes a Prompt-driven Adaptive Token (PrATo) pruning method to selectively reduce the processing of irrelevant tokens in the segmentation pipeline. The prompt-based spatial prior helps to rank the tokens according to their relevance. Tokens with low-relevance scores are down-weighted, ensuring that only the relevant ones are propagated for processing across subsequent stages. This data-driven pruning strategy improves segmentation accuracy and inference speed by allocating computational resources to essential regions. The proposed framework is integrated with several state-of-the-art models to facilitate the elimination of irrelevant tokens, thereby enhancing computational efficiency while preserving segmentation accuracy. The experimental results show a reduction of $\sim$ 35-55% tokens; thus reducing the computational costs relative to baselines. Cost-effective medical image processing, using our framework, facilitates real-time diagnosis by expanding its applicability in resource-constrained environments.
Index Terms:
@@ I Introduction
Precise segmentation of anatomical or pathological areas in medical images is essential for diagnosis, treatment planning, surgical navigation, and personalized medicine. It enables clinicians to measure disease burden, assess treatment response, and improve patient outcomes. Deep learning [1] has revolutionized medical image analysis by offering powerful algorithms to automate complex tasks such as segmentation. Its ability to learn intricate patterns from large datasets without much human intervention has made it a popular choice. Vision Transformers (ViTs) [2] have recently gained prominence in the development of algorithms for computer vision tasks, including medical image segmentation [3].
@@
ViTs employ a self-attention mechanism to capture long-range dependencies in images, thus facilitating global context awareness. This is beneficial for segmenting anatomical structures that exhibit various shapes and sizes. The global view enables ViTs to model interactions between far-apart image regions, which can be difficult for convolutional networks with their restricted receptive fields [4]. Several ViT-based segmentation approaches viz., UNETR [5], TransUNET [6], Swin-UNETR [7] and SegFormer [8] have reported encouraging performance in medical image segmentation, achieving state-of-the-art results across various benchmarks.
@@
Despite impressive performance, ViTs face limitations in resource-constrained environments due to the quadratic computational complexity of self-attention. This scales poorly with increasing input image dimensions. Consequently, the deployment of ViT-based segmentation models in real-time clinical environments becomes challenging, limiting their applicability in practice.
Model compression techniques, such as pruning, are an effective way to mitigate the computational expenses associated with ViTs. Pruning reduces the computational cost of a model by removing less significant elements from its architecture [9]. Many tokens processed by ViTs are semantically redundant [10]. Token pruning increases efficiency by reducing the number of processed tokens.
Static pruning [11, 12, 13] removes the fixed components of the model architecture, before inference, irreversibly eliminating the components considered less significant. For example, TRAM [14] modeled the self-attention mechanism through an MLP layer, with tokens corresponding to the input nodes of the MLP. The association between different tokens was represented by the weights of the MLP. The relevance score of each token was computed recursively on the MLP representation of the self-attention mechanism. The tokens with the lowest scores were discarded in the early layers to ensure that only relevant tokens were forwarded across the network. Intrahead pruning [15] eliminated certain rows of the weight matrices to produce query, key, and value vectors for the self-attention mechanism. An importance score was calculated for each row to eliminate irrelevant entries. A supplementary relationship matrix was preserved to address dependency conflicts with subsequent layers. Shunted Self-Attention [16] performed token reduction by merging across different attention heads to efficiently learn multiscale representations.
@@
Although effective in minimizing model size and computational expense, static pruning is less flexible [17]. Its rigid nature limits the ability to adjust to the unique features of each input image. This constraint poses significant challenges for the analysis of medical images, where the variability between images can be considerable.
In contrast, dynamic pruning [18] adaptively adjusts the computation inference by selectively deactivating model components based on input. Methods like DynamicViT [19],[20] and [21] incorporated lightweight modules to predict the relevance scores of tokens at each stage of the network. The irrelevant tokens with respect to the input image were discarded. Evo-Vit [22] dynamically identified tokens, as informative or placeholder, based on a global class attention score. The placeholder tokens represented irrelevant ones, which were retained simply to maintain the grid structure of the attention map. They were aggregated into a single representative value which was then processed by the self-attention mechanism to reduce computations.
Other works focused on architectural modifications of ViTs to reduce the total number of tokens. TA-ASF [23] implemented a two-stage token pruning method. An importance score was assigned to each token, followed by dividing the entire set of tokens into high- and low-importance sets. A subset of tokens was sampled from both groups to preserve low-importance tokens that might be of global importance in the later layers of the network. Subsequently, the tokens in the subset were merged on the basis of their similarity. The soft top-K token pruning [24] generated a score for each token with a lightweight module. Random noise was added to the scores to prevent the same tokens from being chosen repeatedly. The top K score tokens were subsequently selected in the forward pass. The algorithm next applied a differentiable function to generate probabilities of selecting a token, which were then used to update the score prediction network through backpropagation.
Early exit strategies [25, 26] save computation by bypassing the processing of specific layers for a subset of tokens. However, they might introduce representation inconsistencies within the network. The expansion of the tokens [27] improved the training time of ViT by gradually growing the total number of tokens from a small set of initial seed tokens. Each stage introduced new tokens different from the existing ones. The residual tokens were aggregated and processed along with the selected set.
Although existing pruning techniques have significantly increased the efficiency of ViTs, there exist several drawbacks. Methods performing unstructured pruning hamper the grid-structure of the tokens; thereby, making them incompatible for hierarchical ViT models like Swin UNETR and UNETR. Hard token pruning strategies pose difficulty due to the loss of essential fine-grained information required to reconstruct the segmentation output. Incorporating auxiliary modules to calculate the token relevance scores increases the complexity and parameter count of the model. The importance scores for weighting tokens in dynamic token pruning approaches [19, 20, 21] are often learned without explicit consideration of prior spatial knowledge about the varying target anatomical structures. This might lead to suboptimal segmentation results.
@@
To address these limitations, we introduce a Prompt-driven Adaptive Token (PrATo) pruning approach that explicitly incorporates structured spatial priors into the pruning process. In contrast to techniques that depend exclusively on auxiliary modules or heuristics for selecting tokens, our method utilizes the spatial information embedded in the box prompts to direct token retention. It facilitates input-specific adaptation of the network computation. This guarantees the preservation of tokens essential for depicting the structure of the target object, while redundant or less informative tokens are removed. By directly incorporating this spatial constraint, the proposed method improves the precision of the segmentation and optimizes computational efficiency. The research contributions are summarized as follows.
  • A novel prompt-driven dynamic token pruning framework to increase the efficiency of ViTs by integrating auxiliary spatial prompts. This helps to localize relevant regions related to the target structure.
  • An efficient entropy-based scoring mechanism to quantify the relevance of different image tokens. This parameter-free approach prevents additional computational overhead during training and inference.
  • The generalizability of the method is demonstrated across different state-of-the-art ViT-based medical image segmentation models on different publicly available datasets.
@@
Section II provides a comprehensive overview of the steps involved in the PrATo token pruning framework. The experimental results and details of implementation are presented in Section III, in the publicly available datasets ACDC [28] and ISIC [29]. Qualitative and quantitative results derived from the application of our framework, as embedded in state-of-the-art ViT-based segmentation models, are described. Section V provides a summary of the concluding remarks.
II Methodology
This section details our proposed PrATo framework for integration into ViT blocks of the prevalent ViT-based medical image segmentation models. Recent high-performance medical image segmentation models employ an U-shaped encoder-decoder architecture [30]. PrATo can be seamlessly integrated at different stages of segmentation models to adaptively remove irrelevant tokens from being processed in subsequent steps of the network.
II-A Token generation in ViT
@@
ViTs transform an input image (or feature map volume) $I\in\mathbb{R}^{C\times H\times W}$ into a sequence of tokens [2]. Here, $H$, $W$ and $C$ represent the height, width, and channel dimensions of $I$. The input is partitioned into a set of non-overlapping patches $Z;|Z|=\frac{HW}{p^{2}}$, with each patch having dimension $p^{2}$. They are flattened and linearly projected in $C^{\prime}$-dimensional embedding space, resulting in the sequence $P^{\prime}\in\mathbb{R}^{Z\times C^{\prime}}$. These embedded vectors are called tokens, with $Z$ denoting the total number of tokens generated. Positional embedding is added to the tokens to retain their spatial locations. $P_{0}\in\mathbb{R}^{Z\times C^{\prime}}$ is the sequence of position-sensitive tokens that serve as input to a transformer encoder.
@@
This encoder block consists of alternating layers of multihead self-attention units ($\mu$) and feed-forward network ($\Phi$) units. The $\mu$ captures global contextual relationships between different tokens, performing self-attention parallelly across multiple heads, to operate in various representational subspaces. $\Phi$ is independently applied to the output of each head, to enhance their representational capacity by including an additional non-linear transformation. The operations within a transformer block are expressed as:
$Y_{out}=\Phi[LN\{\mu(LN(P_{0}))\}]+P_{0}$ (1)
Here, $LN$ [31] is the Layer Normalization and $Y_{out}\in\mathbb{R}^{Z\times C^{\prime}}$ is the final output of the transformer block.
II-B Prompt-driven adaptive pruning
@@
Prompts serve as a prior by offering high-level, task-specific information [32]. This directs the model to focus on segmenting specific structures within the input image, resulting in effective segmentation. Consequently, it allows token selection based on contextual information. This framework is referred to as prompt-driven adaptive pruning, where essential tokens are preserved, as shown in Fig. 1.
Figure 1: Illustration of the proposed PrATo framework for pruning ViT tokens.
Figure 2: Illustration of the computation of pruning mask $\beta$ from the similarity score map $S$.
$Y_{out}$ is spatially rearranged to 2D feature maps of dimension $\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$, where $H^{\prime}\times W^{\prime}=Z$. This reshaping maintains spatial coherence as the token order in $Y_{out}$ corresponds to the original spatial locations of the corresponding image patches. Subsequently, it is fed to RoIAlign [33] along with box-prompts to extract region-specific features $F\in\mathbb{R}^{k\times k\times C^{\prime}}$, as shown in Fig. 1. The box prompts serve as spatial priors, highlighting the most informative regions within the input. RoIAlign performs a soft crop and extracts the region-specific feature embeddings corresponding to the Region-of-Interest (RoI) defined in the prompt from $Y_{out}$. Each token in $Y_{out}$ is a contextual embedding of an image patch. Consequently, the ViT token grid will misalign with the box-prompt coordinates, making naive selection of tokens by hard-cropping infeasible. RoIAlign ensures a spatially coherent feature representation, unlike naive selection of tokens to preserve spatial consistency.
@@
$F$ is a summarization of the prompted region generated by RoIAlign. However, it is unable to identify globally relevant ViT tokens alone to segment the target structures. Therefore, a similarity score map $S\in\mathbb{R}^{M\times Z}$ is calculated between the region-specific feature, encoded by $F$ and $Y_{out}$, to measure the relevance of each token. The $F$ and $Y_{out}$ are first transformed into lower-dimensional embeddings $V_{1}\in\mathbb{R}^{M\times d_{V}}$ and $V_{2}\in\mathbb{R}^{Z\times d_{V}}$, respectively, with projection matrices $f_{1}\in\mathbb{R}^{C^{\prime}\times d_{V}}$ and $f_{2}\in\mathbb{R}^{C^{\prime}\times d_{V}}$. This standardizes the feature representations by mapping them from a specific architectural dimension $C^{\prime}$ to a consistent low-dimensional embedding space $d_{V}$. $M$ denotes the number of feature vectors in $V_{1}$. The similarity score is computed as
$S(V_{1},V_{2})=V_{1}V_{2}^{T}/\sqrt{d_{V}}.$ (2)
@@
The next step involves computing the prune mask $\beta\in\mathbb{R}^{Z\times C^{\prime}}$ from the similarity score mask $S$ to filter out irrelevant tokens, as shown in Fig. 2. $S$ is normalized along the token dimension to generate a probability distribution $\rho\in\mathbb{R}^{M\times Z}$ by softmax operation. The probability $\rho_{ij}$ of the $i$ th feature vector of $V_{1}$ ($V_{1i}$) that it is relevant to the $j$th feature vector of $V_{2}$ ($V_{2j}$) is given as $\rho_{ij}=\frac{e^{(S_{ij})}}{\sum_{k=1}^{Z}e^{S_{ik}}}$. $S_{ij}$ denotes the similarity score between $V_{1i}$ and $V_{2j}$. The softmax amplifies high similarities while suppressing lower ones, to distinctly recognize confident matches. Next, the Shannon entropy [34] $H(.)$ is calculated for the probability distribution of each $V_{1i}$, over all ViT tokens. This yields
@@
[…]
An inverse entropy weighting scheme is applied to prioritize tokens having lower uncertainty values. The entropy values are first ranked in ascending order. The ranks $R_{i}|i\in[1,M]$ are normalized to ensure their uniform spacing between 0 and 1. Inverse entropy weights $\Omega_{i}=1-R_{i}$, with $\Omega=\{\Omega_{i}|i\in[1,M]\}$, ensure that higher entropy values (more uncertainty) are assigned to lower weights and vice versa.
Lemma 1 (Uniformity of Weights).
@@
The weighted similarity score map becomes $\tilde{S}=S.\Omega$, the overall relevance $r_{i}$ of the $i$ th symbol being the mean of the weighted similarity scores.
@@
The final step modifies the original tokens based on their relevance scores. A binary mask $\beta\in\{0,1\}^{Z}$ is generated by applying a threshold $\tau$ on the relevance scores as
@@
Here, $\sigma$ is the sigmoid function to squash relevance scores in the range $[0,1]$, and $\beta$ is applied to ViT tokens $Y_{out}$ to effectively remove tokens that are deemed irrelevant based on our proposed prompt-guided weighting and thresholding. The masked token $Y_{out}^{\prime}$ is generated by element-wise multiplication ($\odot$) of $\beta$ with $Y_{out}$, i.e. $Y_{out}^{\prime}=\beta\odot Y_{out}.$
Lemma 2 (Retaining high-relevance token set).
If $H(V_{1i})$ is low, there exists a typical set of high-relevance tokens $S_{T}$ which concentrates the effective information capacity. For the threshold value $\tau$, $S_{T}\subseteq S_{P}$, where $S_{P}$ is the final set of retained tokens.
Proof.
Since $H(V_{1i})$ is low, it satisfies $H(V_{1i})\leq\log_{2}(Z)-\Delta$, for some token reduction factor $\Delta>0$.
By fundamental property of typical sets,
$\exists$ $S_{T}\subset Z$ such that: $S_{T}\leq Z.2^{-\Delta}$ and $S_{T}$ concentrates the effective information capacity as follows:
$\sum_{k\in S_{T}}\rho_{ik}\geq 1-\epsilon;for\>\epsilon>0$